OpenStudy (anonymous):

CORRECT MY ANSWER IF ITS WRONG At the state fair, admission at the gate is $10 . In addition, the cost of each ride is $4 . Suppose that Hans will go on x rides. Hans wants the total number of dollars he spends on admission and rides to be at most t . Using the values and variables given, write an inequality describing this.

OpenStudy (anonymous):

10x+4x>=14

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

I think it should be 10+4x=t

OpenStudy (anonymous):

no greater than symbols?

OpenStudy (da_scienceman):

10+4x<=t?maybe?

OpenStudy (anonymous):

I don't think so because if you plug in 5 for x then it still be 30 on both sides so it cant be less then or greater than t

OpenStudy (da_scienceman):

what is d value of t apriori?

OpenStudy (anonymous):

you sure scienceman?

OpenStudy (da_scienceman):

no u dont..so total number of rides by hans=4x, gate fee =10dollars, maximum amount = t dollars. id total spent muzst be at most t dollars, then total spent by hans = 10 dollars for gate fee + 4x for rides. hence if the maximum bound is t, then inequality is given by: 10+4x <=t! u say its wrong?

OpenStudy (anonymous):

its correct scienceman

OpenStudy (anonymous):

THANKS

OpenStudy (da_scienceman):

sure slimshadyjim! i only give ideas but had to work this one to clear d doubts...no worries.

OpenStudy (da_scienceman):

ok slimjim no shady sry°